OpenStudy (anonymous):

What value of x satisfies cot(90°−x)= –√3/3?

OpenStudy (solomonzelman):

cot of what equals -3√3 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes that was great answer~

OpenStudy (anonymous):

Please Help mine~~~

OpenStudy (solomonzelman):

cot v = - √3/3 --> tan v = -3 / √3 = -√3

OpenStudy (anonymous):

the question ask for it degrees??

OpenStudy (solomonzelman):

I mean tan^(-1) -√3

rishavraj (rishavraj):

apply cot(90-x) = tan x

OpenStudy (solomonzelman):

Basically, cot(90-x)= -√3 / 3 tan(90-x) = -√3 tan^(-1) -√3 = -60º (90-x)º = 60º

rishavraj (rishavraj):

and then rationalise or directly divide the RHS

OpenStudy (solomonzelman):

x=30

OpenStudy (anonymous):

these are the answer choices 120° 240° 210° 150°

OpenStudy (anonymous):

did i typed the question wrong?

rishavraj (rishavraj):

no i dont think so?.... bt why did u ask tht

OpenStudy (anonymous):

just making sure because the answer solomonzelman gave me is not a answer choice

rishavraj (rishavraj):

the answer would be 120...

Parth (parthkohli):

\[\cot (90 - x) = \tan(x)\]

OpenStudy (gudden):

The question you posted is not wrong David.. It's just that the values reoccur... :)

OpenStudy (anonymous):

thanks you guys

OpenStudy (anonymous):

thanks